OpenStudy (anonymous):

Change each fraction to a whole number or a mixed number. 4/3 =

OpenStudy (anonymous):

3/3=1 and 4/3-3/3=1/3 so its 1 1/3

OpenStudy (anonymous):

im stupid........bt i dont get it still

OpenStudy (anonymous):

well 3/3 is a whole

OpenStudy (anonymous):

when adding or subtracting you need the same bottom number

OpenStudy (anonymous):

the the top numbers you just add or substract like you normally would

OpenStudy (anonymous):

and then finally you'd put the bottom number that was the same/common to them both

OpenStudy (anonymous):

ok? understand? if not i'll try and explain it again but ur gonna have to tell me what u dont understand tho

OpenStudy (anonymous):

I DONT UNDERSTAND DA PART WEN U MULTIPLY....WAT DO U MULTIPLY?

OpenStudy (anonymous):

u multiply it so they have the same bottom number

OpenStudy (anonymous):

e.g. the previous q's where you had 12 and 3 as ur bottom number you do 3X4 to get 12 like the other bottom number u see?

OpenStudy (anonymous):

SO I MULTIPLY 4/3 <-

OpenStudy (anonymous):

no beause ur not adding, or ru?

OpenStudy (anonymous):

ur qs here says put it with a whole number

OpenStudy (anonymous):

OK AMA DO A QUESTION AN U TELL ME IF IM WRONG OR RITE IS DAT COOL?

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

sure go ahead

OpenStudy (anonymous):

15/4 SO IM GOIN TO DIVIDE 4/15 AN I GET 3 SUBTRACT 15-12 AN I GET 3 SO IS THE ANSWER 3 3/4?

OpenStudy (anonymous):

YAY! U GOT IT :)

OpenStudy (anonymous):

U DA BEST ...........!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

NO UR UR THE ONE WHO DID IT!! CONGRATS :)

OpenStudy (anonymous):

LOL TANK U MA

OpenStudy (anonymous):

XD LOL

OpenStudy (anonymous):

can u tell me how old ur? just wonderin, dont have to tell me if u dont wanna

OpenStudy (anonymous):

19 i kno im pretty stupid

OpenStudy (anonymous):

an u ???

OpenStudy (anonymous):

17 soz but i thought you were like 13 :P rly 19 and this was ur hw?

OpenStudy (anonymous):

if u ever need help, im happy to help :)

OpenStudy (anonymous):

if u want write me a message at ysanchez@hom.rescare.com

OpenStudy (anonymous):

what now?

OpenStudy (anonymous):

wat u mean?

OpenStudy (anonymous):

dnt worry

OpenStudy (anonymous):

DONE!

OpenStudy (anonymous):

YOU DID IT????

OpenStudy (anonymous):

yerr, i send u a message but what is it coz its not hotmail, gmail or yahoo so what is that, i ment ur email address thing incase u dont kno what im on

OpenStudy (anonymous):

UM IS MY SCHOOL ONLINE TING IDK DEY JST GAVE US DAT SO I USE DAT CUZ I FORGOT MY YAHOO

OpenStudy (anonymous):

lol, k then

OpenStudy (anonymous):

DID U GET MA REPLY ON UR EMAIL

OpenStudy (anonymous):

yerr

OpenStudy (anonymous):

SORRY.....

OpenStudy (anonymous):

??? for???

OpenStudy (anonymous):

BOUT DA MESSAGE ON EMAIL

OpenStudy (anonymous):

dnt be :)

OpenStudy (anonymous):

i mean ur just bein nice

OpenStudy (anonymous):

.......YEA